# Mechanistic Fine-tuning for In-context Learning

### v1

@@ Abstract
In-context Learning (ICL) utilizes structured demonstration-query inputs to induce few-shot learning on Language Models (LMs), which are not originally pre-trained on ICL-style data. To bridge the gap between ICL and pre-training, some approaches fine-tune LMs on large ICL-style datasets by an end-to-end paradigm with massive computational costs. To reduce such costs, in this paper, we propose Attention Behavior Fine-Tuning (ABFT), utilizing the previous findings on the inner mechanism of ICL, building training objectives on the attention scores instead of the final outputs, to force the attention scores to focus on the correct label tokens presented in the context and mitigate attention scores from the wrong label tokens. Our experiments on 9 modern LMs and 8 datasets empirically find that ABFT outperforms in performance, robustness, unbiasedness, and efficiency, with only around 0.01% data cost compared to the previous methods. Moreover, our subsequent analysis finds that the end-to-end training objective contains the ABFT objective, suggesting the implicit bias of ICL-style data to the emergence of induction heads. Our work demonstrates the possibility of controlling specific module sequences within LMs to improve their behavior, opening up the future application of mechanistic interpretability.
@@
[…]
@@ 1 Introduction
Our experiments on 9 modern (L)LMs and 8 downstream datasets demonstrate that ABFT significantly improves ICL performance with satisfactory efficiency, robustness, unbiasedness, and harmlessness, even outperforming previous works of end-to-end fine-tuning the whole model on massive datasets that are approximately larger than ours. Moreover, our analysis finds that the ABFT objective is an implicit bias of direct end-to-end training objective on ICL-style data, suggesting that the causal language modeling on the ICL-style data may naturally evoke the emergence of induction heads.
@@ 3 Attention Behavior Fine-tuning
Globally, ABFT utilizes such a pipeline: (1) Dataset Building: from a selected downstream dataset, we build a training set composed of ICL-style sequences as shown in Fig. 1 (A). (2) Feed-forward Calculation: For each training sample, as shown in Fig. 1 (B), we conduct a standard feed-forward calculation on the pre-trained LM, and collect the attention matrices of all the attention heads in all the layers. (3) Loss Calculation: For each attention matrix, we only focus on the attention scores of the last token (i.e., the last row of the attention matrix), where the predictions of queries are made. As shown in Fig. 1 (D), we first filter (detailed below) the non-induction head out, and return a loss of for these heads. Then, for the remaining induction heads, we calculate a loss function composed of a of attention scores on wrong labels and a of attention scores on correct labels (detailed below). (4) Back Propagation: We back-propagate the calculated loss only to the and matrices of every attention head, and update the model parameters.
@@
As shown in Fig. 1 (C), we skip the attention matrices where the attention scores of the last token do not dominate on the label tokens. To identify the attention matrices to skip, in detail, given an attention matrix , where the is the input token sequence length, as mentioned before, we focus on the last row . Given the position index of label tokens as , we calculate the attention score sum on these label tokens as .111For the case shown in Fig. 1 (C), (0-started), and is the sum of the values of the red-highlighted bar and the blue-highlighted bar. Then, we set a threshold , if , we assert the attention head of score is an induction head, and vice versa. We will verify the necessity and benefits of this induction head filter in §5.2.
Model | Param. | Method | SST2 | MR | FP | SST5 | TREC | SUBJ | TEE | TEH | Average |
GPT2-L | 812M | Vanilla | 56.35 | 61.13 | 42.09 | 29.69 | 35.45 | 49.12 | 39.06 | 54.00 | 45.86 |
MetaICL | 85.94 | 80.96 | 37.30 | 42.09 | 33.98 | 50.49 | 45.41 | 54.20 | 53.80 | ||
PICL | 74.70 | 73.34 | 54.49 | 33.79 | 32.91 | 51.37 | 47.46 | 53.42 | 52.68 | ||
Calibrate | 56.35 | 60.94 | 36.91 | 25.10 | 34.08 | 49.32 | 36.72 | 53.52 | 44.12 | ||
ABFT | 88.181.54 | 85.400.95 | 81.301.69 | 36.843.61 | 50.242.49 | 61.994.39 | 46.513.11 | 55.200.20 | 63.21 | ||
GPT2-XL | 1.61B | Vanilla | 67.87 | 69.53 | 51.07 | 30.66 | 35.25 | 50.98 | 42.68 | 53.03 | 50.13 |
PICL | 74.80 | 74.32 | 51.17 | 32.71 | 33.20 | 51.46 | 47.95 | 53.42 | 52.38 | ||
Calibrate | 68.16 | 75.00 | 36.43 | 28.52 | 35.55 | 50.10 | 39.26 | 51.56 | 48.07 | ||
ABFT | 87.921.47 | 86.521.50 | 87.670.45 | 37.552.67 | 51.832.73 | 75.072.96 | 60.015.38 | 55.350.04 | 67.74 | ||
Falcon3 | 7.46B | Vanilla | 91.11 | 92.77 | 85.35 | 46.00 | 50.00 | 62.60 | 60.55 | 52.05 | 67.55 |
Calibrate | 90.53 | 93.07 | 82.71 | 44.04 | 54.30 | 62.40 | 54.79 | 51.76 | 66.70 | ||
ABFT | 92.140.21 | 92.170.04 | 96.140.36 | 47.320.16 | 75.810.19 | 94.870.82 | 67.970.25 | 70.340.22 | 79.59 | ||
Llama3 | 8.03B | Vanilla | 89.35 | 92.87 | 75.78 | 44.24 | 55.76 | 62.30 | 57.91 | 54.59 | 66.60 |
Calibrate | 90.04 | 93.36 | 43.95 | 41.60 | 54.39 | 65.23 | 54.79 | 52.83 | 62.02 | ||
ABFT | 93.141.39 | 92.500.84 | 94.021.72 | 52.102.31 | 73.091.11 | 92.701.44 | 72.021.81 | 72.025.87 | 80.20 | ||
DeepSeek-R1 Dist. Qwen 4-bit, LoRA | 14.8B | Vanilla | 90.92 | 91.21 | 92.18 | 46.97 | 62.50 | 66.60 | 66.60 | 63.87 | 72.61 |
Calibrate | 90.04 | 91.41 | 92.68 | 46.09 | 61.62 | 65.43 | 65.33 | 62.30 | 71.86 | ||
ABFT | 93.511.22 | 91.850.34 | 91.174.44 | 46.091.27 | 69.142.54 | 92.923.27 | 69.820.00 | 71.193.42 | 78.21 | ||
Qwen2.5 4-bit, LoRA | 32.8B | Vanilla | 93.85 | 94.43 | 86.23 | 47.17 | 58.40 | 87.50 | 65.14 | 69.63 | 75.29 |
Calibrate | 93.75 | 94.82 | 74.22 | 44.82 | 58.79 | 84.08 | 63.96 | 63.96 | 72.30 | ||
ABFT | 94.920.00 | 94.830.10 | 94.040.00 | 48.490.05 | 69.240.10 | 96.000.00 | 69.290.24 | 70.310.00 | 79.64 | ||
SimpleScaling s1.1 4-bit, LoRA | 32.8B | Vanilla | 94.82 | 94.24 | 91.11 | 50.20 | 69.63 | 89.65 | 68.36 | 72.17 | 78.77 |
Calibrate | 94.43 | 93.85 | 88.96 | 48.63 | 68.07 | 89.26 | 68.55 | 72.36 | 78.02 | ||
ABFT | 94.920.10 | 94.290.05 | 96.000.00 | 49.950.05 | 72.460.39 | 95.710.10 | 71.730.05 | 73.100.04 | 81.02 | ||
Llama3 4-bit, LoRA | 43.2B | Vanilla | 93.26 | 94.04 | 73.92 | 49.41 | 58.98 | 71.58 | 62.60 | 66.70 | 71.31 |
Calibrate | 95.02 | 93.07 | 54.20 | 44.53 | 59.08 | 72.56 | 61.03 | 65.82 | 68.16 | ||
ABFT | 95.020.10 | 93.850.10 | 94.870.05 | 48.100.14 | 64.700.14 | 90.090.05 | 69.240.29 | 70.850.15 | 78.34 | ||
Llama3 4-bit, LoRA | 55.6B | Vanilla | 93.94 | 92.19 | 78.81 | 51.37 | 67.19 | 66.70 | 56.44 | 60.94 | 70.95 |
Calibrate | 92.29 | 92.77 | 69.92 | 50.49 | 68.75 | 65.92 | 58.11 | 62.70 | 70.12 | ||
@@
As shown in Fig. 1 (D), given an attention matrix , if judged as a non-induction head by the aforementioned head filter, the loss for is assigned to . Else, we conduct the following calculation: given the position index of label tokens consistent with the ground-truth label of the query as , and the others ,222For the case shown in Fig. 1 (C), (the position of “positive”), (the position of “negative”). we calculate the loss from the last row () of as:
That is, as shown in Fig. 1 (D), we the “wrong” attention scores towards the label tokens different from the query’s ground-truth with magnitude , and the “correct” attention scores with magnitude . These two terms in the loss function may seem redundant, but we will demonstrate in §5.3 that they actually contain antagonistic implicit biases, therefore, the factors and should be balanced well.
@@ 4.2 Main Results
Model | Method | Param.* | Time | Acc | Acc |
Llama3 8.03B 4-bit, LoRA | Vanilla | - | - | 66.02 | |
E2E FT | 0.5B | 78.33 | 61.74 | ||
ABFT | 6.8M | 72.54 | 64.34 | ||
DeepSeek-R1 14.8B 4-bit, LoRA | Vanilla | - | - | 72.61 | |
E2E FT | 0.8B | 78.26 | 63.62 | ||
ABFT | 12M | 78.21 | 67.21 | ||
Qwen2.5 32.8B 4-bit, LoRA | Vanilla | - | - | 75.29 | |
E2E FT | 0.8B | 82.09 | 62.24 | ||
ABFT | 17M | 79.64 | 64.96 | ||
Llama3 55.6B 4-bit, LoRA | Vanilla | - | - | 70.95 | |
E2E FT | 1.1B | 82.80 | 64.86 | ||
ABFT | 33M | 79.52 | 67.32 |
@@
Moreover, a known concern in ICL is the bias toward seen labels when ground-truth labels are absent in demonstrations (i.e., ) Zhao et al. (2021); Cho et al. (2025a), which can lead to incorrect predictions. Our testing on such scenario with and without ABFT in Fig. LABEL:fig:unseen shows that (see Appendix A.4 for experiment details): ABFT mitigates this issue via the , which penalizes incorrect labels during training and reduces the bias effects of induction heads. Notably, ABFT outperforms the 0-shot setting under unseen-label, suggesting the existence or emergence of unknown mechanisms that enable demonstrations in other categories to enhance ICL333Since that: such a phenomenon contrasts with existing views Cho et al. (2025a), where the explicit copying by the induction head is the only channel through which information is transferred from the demonstrations to the query (in Fig. 10, we show that the induction heads in ABFT model are almost fully suppressed in unseen-label scenario), where unseen-label demonstrations are harmful to ICL..
@@ 4.3 Comparison against End-to-end Fine-tuning
We test the accuracy against the training set size as a metric of data efficiency, for both ABFT and E2E fine-tuning, as shown in Fig. LABEL:fig:data_eff (refer to Appendix C.1 for results on other models). In the results, ABFT and E2E fine-tuning consistently benefit from more data samples, and in few-shot scenarios (), ABFT and E2E fine-tuning act equally, while E2E fine-tuning acts better when more training data is given. However, given the low-resource objective of ICL, and also the far more expensive time and memory cost of E2E fine-tuning, we can claim that ABFT has an advantage in the few-shot and low-resource scenario.
@@ 5.2 Ablation Analysis
Model | Method | Time | Acc. |
Falcon3 7.46B | Vanilla | - | 67.55 |
ABFT | 79.59 | ||
w/o PID, | 76.86 | ||
w/o PID, w/o Head Filter | 75.57 | ||
w/o PID, | 72.13 | ||
w/o PID, | 56.47 | ||
Llama3 8.03B | Vanilla | - | 66.60 |
ABFT | 80.20 | ||
w/o PID, | 80.07 | ||
w/o PID, w/o Head Filter | 70.39 | ||
w/o PID, | 63.54 | ||
w/o PID, | 58.79 | ||
DeepSeek-R1 14.8B Dist. Qwen 4-bit, LoRA | Vanilla | - | 72.61 |
ABFT | 78.21 | ||
w/o PID, | 73.35 | ||
w/o PID, w/o Head Filter | 73.51 | ||
w/o PID, | 72.71 | ||
@@ 5.3 Balance the Loss Factor and
and Influence Induction Heads Antagonisticly.
As shown in Table 4, removing either loss term in Eq. 1 degrades performance, indicating that both are essential. Interestingly, the two terms introduce antagonistic implicit biases: the disperses attention across labels, reducing induction heads, while the concentrates attention on specific labels, increasing induction heads through a stepwise positive feedback loop. We track the number of induction heads during training on Llama3 8B and SST2 (see Appendix C.2 for more cases), as shown in Fig. 5, to support this observation.
@@
Ablation studies reveal that too many induction heads hinder fine-tuning and overall performance, as other functional heads are also needed for ICL; whereas an insufficient number prevents the model from handling ICL tasks. To maintain a stable number, we automatically adjust the antagonistic factor in Eq. 1 using a classical PID controller (Appendix A.3)444Since and are antagonistic, controlling alone suffices to stabilize induction head numbers.. As shown in Fig. 5, PID stabilizes induction head count, improves ABFT performance (Table 4), and reduces sensitivity to hyperparameters and , with accuracy remaining stable across settings (Fig. 6).
@@ Conclusion.
The first limitation lies in the fact that the induction head-based explanation of ICL Elhage et al. (2021); Singh et al. (2024); Cho et al. (2025a), so that our proposed ABFT approach, applies only to classification tasks with a finite label set. As mentioned in §4.2 and Fig. LABEL:fig:unseen, since our training objective consists of two factors, our approach is not limited to the simple retrieval setting where the ground-truth label appears in the demonstrations, while extending these methods to open-ended tasks remains an open challenge that requires further investigation on the basic mechanism. Nevertheless, given the current state of research on ICL interpretability, we have made full use of these findings and provided a valuable foundation for advancing model control through the scope of interpretability, i.e., Mechanistic Controllability.

### v2

@@ Abstract
In-context Learning (ICL) utilizes structured demonstration-query inputs to induce few-shot learning on Language Models (LMs), which are not originally pre-trained on ICL-style data. To bridge the gap between ICL and pre-training, some approaches fine-tune LMs on large ICL-style datasets by an end-to-end paradigm with massive computational costs. To reduce such costs, in this paper, we propose Attention Behavior Fine-Tuning (ABFT), utilizing the previous findings on the inner mechanism of ICL, building training objectives on the attention scores instead of the final outputs, to force the attention scores to focus on the correct label tokens presented in the context and mitigate attention scores from the wrong label tokens. Our experiments on 9 modern LMs and 8 datasets empirically find that ABFT outperforms in performance, robustness, unbiasedness, and efficiency, with only around 0.01% data cost compared to the previous methods. Moreover, our subsequent analysis finds that the end-to-end training objective contains the ABFT objective, suggesting the implicit bias of ICL-style data to the emergence of induction heads. Our work demonstrates the possibility of controlling specific module sequences within LMs to improve their behavior, opening up the future application of mechanistic interpretability111Source code of this paper is available at https://github.com/hc495/ICL_head_tuning..
@@
Hakaze Cho Peng Luo2 Mariko Kato1 Rin Kaenbyou2 Naoya Inoue1,3 1Japan Advanced Institute of Science and Technology 2Beijing Institute of Technology3RIKEN Primary Contributor, Correspondence to: yfzhao@jaist.ac.jp
@@ 1 Introduction
Our experiments on 9 modern (L)LMs and 8 downstream datasets demonstrate that ABFT significantly improves ICL performance with satisfactory efficiency, robustness, unbiasedness, and harmlessness, even outperforming previous works of end-to-end fine-tuning the whole model on massive datasets that are approximately larger than ours. Moreover, our analysis shows that the ABFT objective implicitly biases end-to-end training on ICL-style data, where causal language modeling may naturally induce induction heads.
@@ 3 Attention Behavior Fine-tuning
[…]
@@
As shown in Fig. 1 (C), we skip the attention matrices where the attention scores of the last token do not dominate on the label tokens. To identify the attention matrices to skip, in detail, given an attention matrix , where the is the input token sequence length, as mentioned before, we focus on the last row . Given the position index of label tokens as , we calculate the attention score sum on these label tokens as .222For the case shown in Fig. 1 (C), (0-started), and is the sum of the values of the red-highlighted bar and the blue-highlighted bar. Then, we set a threshold , if , we assert the attention head of score is an induction head, and vice versa. We will verify the necessity and benefits of this induction head filter in §5.2.
| Model | Param. | Method | SST2 | MR | FP | SST5 | TREC | SUBJ | TEE | TEH | Average | |||
| GPT2-L | 812M | Vanilla | 56.35 | 61.13 | 42.09 | 29.69 | 35.45 | 49.12 | 39.06 | 54.00 | 45.86 | |||
| MetaICL | 85.94 | 80.96 | 37.30 | 42.09 | 33.98 | 50.49 | 45.41 | 54.20 | 53.80 | |||||
| PICL | 74.70 | 73.34 | 54.49 | 33.79 | 32.91 | 51.37 | 47.46 | 53.42 | 52.68 | |||||
| Calibrate | 56.35 | 60.94 | 36.91 | 25.10 | 34.08 | 49.32 | 36.72 | 53.52 | 44.12 | |||||
| ABFT | 88.181.54 | 85.400.95 | 81.301.69 | 36.843.61 | 50.242.49 | 61.994.39 | 46.513.11 | 55.200.20 | 63.21 | |||||
| GPT2-XL | 1.61B | Vanilla | 67.87 | 69.53 | 51.07 | 30.66 | 35.25 | 50.98 | 42.68 | 53.03 | 50.13 | |||
| PICL | 74.80 | 74.32 | 51.17 | 32.71 | 33.20 | 51.46 | 47.95 | 53.42 | 52.38 | |||||
| Calibrate | 68.16 | 75.00 | 36.43 | 28.52 | 35.55 | 50.10 | 39.26 | 51.56 | 48.07 | |||||
| ABFT | 87.921.47 | 86.521.50 | 87.670.45 | 37.552.67 | 51.832.73 | 75.072.96 | 60.015.38 | 55.350.04 | 67.74 | |||||
| Falcon3 | 7.46B | Vanilla | 91.11 | 92.77 | 85.35 | 46.00 | 50.00 | 62.60 | 60.55 | 52.05 | 67.55 | |||
| Calibrate | 90.53 | 93.07 | 82.71 | 44.04 | 54.30 | 62.40 | 54.79 | 51.76 | 66.70 | |||||
| ABFT | 92.140.21 | 92.170.04 | 96.140.36 | 47.320.16 | 75.810.19 | 94.870.82 | 67.970.25 | 70.340.22 | 79.59 | |||||
| Llama3 | 8.03B | Vanilla | 89.35 | 92.87 | 75.78 | 44.24 | 55.76 | 62.30 | 57.91 | 54.59 | 66.60 | |||
| Calibrate | 90.04 | 93.36 | 43.95 | 41.60 | 54.39 | 65.23 | 54.79 | 52.83 | 62.02 | |||||
| ABFT | 93.141.39 | 92.500.84 | 94.021.72 | 52.102.31 | 73.091.11 | 92.701.44 | 72.021.81 | 72.025.87 | 80.20 | |||||
14.8B | Vanilla | 90.92 | 91.21 | 92.18 | 46.97 | 62.50 | 66.60 | 66.60 | 63.87 | 72.61 | |||
| Calibrate | 90.04 | 91.41 | 92.68 | 46.09 | 61.62 | 65.43 | 65.33 | 62.30 | 71.86 | |||||
| ABFT | 93.511.22 | 91.850.34 | 91.174.44 | 46.091.27 | 69.142.54 | 92.923.27 | 69.820.00 | 71.193.42 | 78.21 | |||||
32.8B | Vanilla | 93.85 | 94.43 | 86.23 | 47.17 | 58.40 | 87.50 | 65.14 | 69.63 | 75.29 | |||
| Calibrate | 93.75 | 94.82 | 74.22 | 44.82 | 58.79 | 84.08 | 63.96 | 63.96 | 72.30 | |||||
| ABFT | 94.920.00 | 94.830.10 | 94.040.00 | 48.490.05 | 69.240.10 | 96.000.00 | 69.290.24 | 70.310.00 | 79.64 | |||||
32.8B | Vanilla | 94.82 | 94.24 | 91.11 | 50.20 | 69.63 | 89.65 | 68.36 | 72.17 | 78.77 | |||
| Calibrate | 94.43 | 93.85 | 88.96 | 48.63 | 68.07 | 89.26 | 68.55 | 72.36 | 78.02 | |||||
| ABFT | 94.920.10 | 94.290.05 | 96.000.00 | 49.950.05 | 72.460.39 | 95.710.10 | 71.730.05 | 73.100.04 | 81.02 | |||||
43.2B | Vanilla | 93.26 | 94.04 | 73.92 | 49.41 | 58.98 | 71.58 | 62.60 | 66.70 | 71.31 | |||
| Calibrate | 95.02 | 93.07 | 54.20 | 44.53 | 59.08 | 72.56 | 61.03 | 65.82 | 68.16 | |||||
| ABFT | 95.020.10 | 93.850.10 | 94.870.05 | 48.100.14 | 64.700.14 | 90.090.05 | 69.240.29 | 70.850.15 | 78.34 | |||||
55.6B | Vanilla | 93.94 | 92.19 | 78.81 | 51.37 | 67.19 | 66.70 | 56.44 | 60.94 | 70.95 | |||
| Calibrate | 92.29 | 92.77 | 69.92 | 50.49 | 68.75 | 65.92 | 58.11 | 62.70 | 70.12 | |||||
@@
As shown in Fig. 1 (D), given an attention matrix , if judged as a non-induction head by the aforementioned head filter, the loss for is assigned to . Else, we conduct the following calculation: given the position index of label tokens consistent with the ground-truth label of the query as , and the others ,333For the case shown in Fig. 1 (C), (the position of “positive”), (the position of “negative”). we calculate the loss from the last row () of as:
[…]
@@ 4.2 Main Results
| Model | Method | Param.* | Time | Acc | Acc | |||
Vanilla | - | - | 66.02 | ||||
| E2E FT | 0.5B | 78.33 | 61.74 | |||||
| ABFT | 6.8M | 72.54 | 64.34 | |||||
Vanilla | - | - | 72.61 | ||||
| E2E FT | 0.8B | 78.26 | 63.62 | |||||
| ABFT | 12M | 78.21 | 67.21 | |||||
Vanilla | - | - | 75.29 | ||||
| E2E FT | 0.8B | 82.09 | 62.24 | |||||
| ABFT | 17M | 79.64 | 64.96 | |||||
Vanilla | - | - | 70.95 | ||||
| E2E FT | 1.1B | 82.80 | 64.86 | |||||
| ABFT | 33M | 79.52 | 67.32 | |||||
@@
[…]
@@ 4.3 Comparison against End-to-end Fine-tuning
We test the accuracy against the training set size as a metric of data efficiency, for both ABFT and E2E fine-tuning, as shown in Fig. 3 (refer to Appendix C.1 for results on other models). In the results, ABFT and E2E fine-tuning consistently benefit from more data samples, and in few-shot scenarios (), ABFT and E2E fine-tuning act equally, while E2E fine-tuning acts better when more training data is given. However, given the low-resource objective of ICL, and also the far more expensive time and memory cost of E2E fine-tuning, we can claim that ABFT has an advantage in the few-shot and low-resource scenario.
@@ 5.2 Ablation Analysis
| Model | Method | Time | Acc. | ||||
Vanilla | - | 67.55 | ||||
| ABFT | 79.59 | ||||||
| w/o PID, | 76.86 | ||||||
| w/o PID, w/o Head Filter | 75.57 | ||||||
| w/o PID, | 72.13 | ||||||
| w/o PID, | 56.47 | ||||||
Vanilla | - | 66.60 | ||||
| ABFT | 80.20 | ||||||
| w/o PID, | 80.07 | ||||||
| w/o PID, w/o Head Filter | 70.39 | ||||||
| w/o PID, | 63.54 | ||||||
| w/o PID, | 58.79 | ||||||
Vanilla | - | 72.61 | ||||
| ABFT | 78.21 | ||||||
| w/o PID, | 73.35 | ||||||
| w/o PID, w/o Head Filter | 73.51 | ||||||
| w/o PID, | 72.71 | ||||||
@@ 5.3 Balance the Loss Factor and
[…]
@@
Ablation studies reveal that too many induction heads hinder fine-tuning and overall performance, as other functional heads are also needed for ICL; whereas an insufficient number prevents the model from handling ICL tasks. To maintain a stable number, we automatically adjust the antagonistic factor in Eq. 1 using a classical PID controller (Appendix A.3)555Since and are antagonistic, controlling alone suffices to stabilize induction head numbers.. As shown in Fig. 5, PID stabilizes induction head count, improves ABFT performance (Table 4), and reduces sensitivity to hyperparameters and , with accuracy remaining stable across settings (Fig. 6).
@@ Conclusion.
The first limitation lies in the fact that the induction head-based explanation of ICL Elhage et al. (2021); Singh et al. (2024); Cho et al. (2025a), so that our proposed ABFT approach, applies only to classification tasks with a finite label set. As mentioned in §4.2 and Fig. 3, since our training objective consists of two factors, our approach is not limited to the simple retrieval setting where the ground-truth label appears in the demonstrations, while extending these methods to open-ended tasks remains an open challenge that requires further investigation on the basic mechanism. Nevertheless, given the current state of research on ICL interpretability, we have made full use of these findings and provided a valuable foundation for advancing model control through the scope of interpretability, i.e., Mechanistic Controllability.